INAF LBC+LUCIFER queue observing, 2010 October 07-08

Observer: V. Testa, K.Boutsia, M. Fumana
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O.Kuhn
Instrument: LBC

Summary

quite cloudy at the beginning of the night
> 2 hrs lost because of mirror problems (Co-pointing did not work and red mirror would not converge)

> 3 hrs lost because of wrongly tilted gratings, diagonal slits, NB4 changed to NB6 and other amenities as such

Data Summary

Project Objects Filter ExpTime Notes
Clementini AndXXI - AndXIX B-bessel, V-bessel 21min + 21min seeing ~1.3'', out of specs
Turatto SN2009ig J, H, Ks 640 s + 896 s + 896 s seeing around 1"
Turatto SN2010hq J, H, Ks 20s + 8s + 8s seeing below 1"
Turatto SN2010hq 200HK 800 s seeing below 1" - Telluric + ARC +FLAT

Details

01:21 - Slew for flats

01:30 - Flats
  • Flat script in binocular mode SkyFlat_BI
  • Flat script in binocular mode SkyFlat_VV
  • Scripts created with Mkskyflat.pl

01:35 - Filter Error (Timeout error Stop) - restart Flat_VV

01:39 - Switch OFF/ON LBC

01:48 - Slew to Pointing Star - ACT0219
  • ptautoadj + mirradjust

01:58 - Problems with the red primary mirror (collimation position error)

02:04 - Slew to standard SA113_BV_VI - dofpia * Hit limits on primary mirror * dofpia, /BackOut

02:20 - Slew to AndXXI (Clementini)
  • Filter system stopped (RED)
  • we have to start from scratch - including pointing correction

02:22 - Slew to pointing star - ACT0338
  • DID NOT SLEW
  • Slew again - correct pointing

02:40 - Hit limits - correct pointing and mirror adjust
  • Co-pointing does NOT work,
  • We decide to go on without co-pointing

02:54 - Slew to AndXXI (Clementini) - dofpia
  • Z7=+2000 (red, given manually)
  • red does not converge

03:21 - STOP - Clear optics - dofpia
  • Olga calls Andrew because red mirror keeps giving coma

03:48 - dofpia
  • z7=+2500, z8=+1000 (red, given manually)
  • red does not converge

03:53 - dofpia, /Redonly
  • interrupted - red does not converge

04:10 - dofpia
  • we try to set values manually at the beginning
  • Z11=+1500 (Blue) and Z7=-2000 Z8=+1500 Z11=+1000 (Red)
  • CONVERGED!!!

04:18 - Exposure - M31_AndXXI_420s_1 (Clementini)
  • seeing: 1.2'' (Red) - >1.3'' (Blue)

04:46 - Slew to AndXIX (Clementini) - dofpia
  • dofpia oscillates a lot, probably because of bad seeing
  • we stop it and try to do the exposure anyway

04:50 - Exposure - M31_AndXIX_420s_1 (Clementini)
  • seeing: ~1.2'' (Red) - ~1.4'' (Blue)
  • seeing out of spec in both OBs

05:18 - SWITCH TO LUCIFER

05:59 - Ready to go, pointing already checked. Acquire Longhetti telluric 1 (Hip13917)

06:06 - Collimated and guiding, acquiring 0001-0004
  • offsetting the telescope seems no to work, both with the N,W,E,S button then x,y shift, change coordinate system RADEC-->DETXY

06:16 - Stop and restart acquisition 0005-0008

06:23 - Telluric1 (Hip13917) 0009-0016

06:29 - After struggling a while to put the object in the slit, we discovered the slit placement in the cabinet has changed wrt the manual and we were using 2" slit instead of 1"
  • have to redo the telluric.
  • Another hour wasted, this was in the wiki but not in the manual;

06:57 - Telluric 1 - Acquisition with the right slit 0029-0030

07:07 - Telluric 1 - Spectra 0031-0038
  • 0031 - don't see the spectrum, only the very right of the chip.....

07:20 - Called Nancy for help
  • it turns out that the 150Ks grating is not in the correct position. So we'll have to do the acquisition, take out instrument setup, then call Nancy who will setup the proper tilt for us.

07:45 - Redo Pointing Correction
  • The slit is not perfectly aligned along Y so likely the target will fall also out of the slit like the telluric does when it is offset by -50" * Tests to see slit position and other technical frames 0039-0064

08:17 - move to another target but before TO needs some minutes to take the enclosure out of the wind (that blows at ~14 m/s)

08:23 - Slew to SN2009ig -acquisition (Turatto) 0065-0066

08:27 - Science script for imaging SN2009ig
  • J 0067-0074, H 0075-0090, K 0091-0106

09:36 - Give telescope to TO for anti-wind positioning

09:47 - With assistance from Nancy we also reposition correctly the 3.75 camera that shows an unusual border.

10:10 - Acquisition SN2010hq imaging (Turatto)
  • harder to recognize the field 0114-0115

10:20 - Science script for imaging SN2010hq
  • J 0116-0119= ; H ==0120-0123 ; Ks 0124-0127

10:26 - SN2010hq Spec Acquisition
  • quite difficult to center object in the slit without a finding chart plus wrong slit (NB4 is now NB6)

10:50 - SN2010hq Spec Science 0131-0136
  • slipped out of the slit, useless
  • need to correct script - repeat acquisition and see if object within the slit 0137-0145

11:44 - reacquired and taken spectra for SN2010hq
  • this time on-slit 0146-0157

12:00 - calibration unit does not come IN, cannot do arcs and flats, Olga tries to insert it by hand.....
  • give up do the telluric

12:15 - Hip34407 - Telluric for SN2010hq
  • telluric acquisition 0158
  • telluric science script (5sec) 0159-0170
  • repeat telluric with higher exposure time (10sec) because flux too low 0171-0182

12:35 - Calibration: Arc and Flats (cal unit activated by hand)
  • FLAT : lamp OFF 0183-0192, lamp ON 0193-202
  • ARC: no arcs ???

13:05 - Darks

That's it, a very painful night

-- VincenzoTesta - 08 Oct 2010
Topic revision: r7 - 11 Oct 2010, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback